# On the Low-Temperature MCMC threshold:

### v2

@@ 12. Deferred proofs
Theorem 12.2.
[Theorem 5.2 of [BDDW08]] Let $n,k,p\in\mathbb{N}$ with $k\leq p$. Suppose $Z\in\mathbb{R}^{n\times p}$ has i.i.d. standard Gaussian entries. Then for every $\delta>0$, there exists a constant $C=C_{\delta}>0$ such that if $n\geq Ck\log\frac{p}{k}$, then $Z$ satisfies the $k$-RIP with restricted isometric constant $\delta$ w.h.p. as $k\rightarrow\infty$.
@@
Finally, we will need two results from [GZ22b] concerning D.L.M.s
Proposition 12.5.
[Proposition F.7 of [GZ22b]] Let $n,p,k\in\mathbb{N}$ with $k\leq\frac{1}{3}p$. Suppose that $X\in\mathbb{R}^{n\times p}$ satisfies the $3k$-RIP with restricted isometry constant $\delta_{3k}<\frac{1}{12}$. Then there is no $\frac{1}{4}$-D.L.M. triplet $(a,b,c)$ with respect to the matrix $X$ with $\lVert a\rVert_{2}^{2}+\lVert b\rVert_{2}^{2}\geq\frac{1}{4}\lVert c\rVert_{2}^{2}$.
Proposition 12.6.
[Proposition F.8 of [GZ22b]] Let $n,p,k\in\mathbb{N}$ with $k\leq\frac{1}{3}p$. Suppose that $X\in\mathbb{R}^{n\times p}$ has i.i.d. $\mathcal{N}(0,1)$ entries. There exist constants $c_{1},C_{1}>0$ such that if $n\geq C_{1}k\log(p/k)$ then w.h.p., there is no $\frac{1}{4}$-D.L.M. triplet $(a,b,c)$ with respect to some sets $\emptyset\neq S_{1},S_{2},S_{3}\subset[p]$ and the matrix $X$ such that the following conditions are satisfied.
@@
    $\lVert a\rVert_{2}^{2}+\lVert b\rVert_{2}^{2}+\lVert c\rVert_{2}^{2}\leq c_{1}\min\left\{\frac{\log p}{\log\log p},k\right\}$
Note that the original proof of Proposition 12.6 in [GZ22b] assumes $n\geq C_{1}k\log(p)$, but the proof is identical if we assume $n\geq C_{1}k\log(p/k)$. Now we have everything we need to prove Lemma 8.4.
@@
Choose $C>0$ large enough so that $X^{\prime}$ satisfies $3k$-RIP (Restricted Isometry Property) with $\delta_{3k}<\frac{1}{12}$ (we know we can do this by Theorem 12.2). Let $T$ denote the support of $v$, and without loss of generality, assume that the support of $v^{*}$ is $[k]$. Denote $m=|T\setminus[k]|=|[k]\setminus T|$.
@@
$\lVert Y-Xv^{\prime}\rVert^{2}_{2}\geq\lVert Y-Xv\rVert^{2}_{2}-\frac{1}{4}n$ (12.1)
@@
$\lVert Xv^{*}+W-Xv+X_{j}-X_{i}\rVert_{2}^{2}\geq\lVert Xv^{*}+W-Xv\rVert^{2}_{2}-\frac{1}{4}n$ (12.2)
@@
Observe that we can rewrite the inequality (12.2) above in terms of $X^{\prime}$ to get that for all $i\in[p]$ and for all $j\in T$, we have
$\lVert X^{\prime}a+X^{\prime}b+X^{\prime}c-a_{i}X^{\prime}_{i}-b_{j}X^{\prime}_{j}\rVert^{2}_{2}\geq\lVert X^{\prime}(a+b+c)\rVert_{2}^{2}-\frac{1}{4}n$ (12.3)
This therefore gives that $(a,b,c)$ is $\frac{1}{4}$-DLM with respect to $[k]\setminus T$, $T\setminus[k]$, $\{p+1\}$, and the matrix X’. Therefore since $X^{\prime}$ satisfies $3k$-RIP with $\delta_{3k}<\frac{1}{12}$, by Proposition 12.5, we have $\lVert a\rVert_{2}^{2}+\lVert b\rVert_{2}^{2}\leq\frac{1}{4}\lVert c\rVert^{2}_{2}$, or equivalently
@@
We therefore apply Proposition 12.6 to conclude that it must be the case that for some positive constant $c_{1}$, with high probability
$c_{1}\min\left\{\frac{\log p}{\log\log p},k\right\}\leq|\text{supp}(a)|+|\text{supp}(b)|+\sigma^{2}\leq\frac{5}{4}\sigma^{2}$
@@
$\frac{4}{5}c_{1}\min\left\{\frac{\log p}{\log\log p},k\right\}\leq\sigma^{2}$
We choose our constant $c>0$ so that $c<\frac{4}{5}c_{1}$, so that $\sigma^{2}<\frac{4}{5}c_{1}\min\{\frac{\log p}{\log\log p},k\}$, and we get a contradiction. ∎
12.2. Proof of Theorem 10.1
Here we prove Theorem 10.1. The strategy here is that we will first prove a similar result for a pure noise model (i.e. with no signal). We then show how the linear model can be reduced to the pure noise model. We again follow [GZ22a]. Note that, as we mentioned also above, the first part of Theorem 10.1 is unchanged from [GZ22a], so we only prove the second part.
12.2.1. The Pure Noise Model
@@
Here there is no “true” vector $v^{*}$; we are simply trying to find the exactly $k$-sparse binary vector that minimises this squared error. Denote the optimal value of this optimization problem by $\psi_{2}$. Theorem 12.7 is our main result for this subsection. It improves Theorem A.1 of [GZ22b] by reducing the assumption that $\Omega(k\log k)=n=o(k\log p)$ (and hence $k=p^{o(1)}$) to $k=o(p^{1/3-\eta})$ for some $\eta>0$ and no assumption on $n$.
Theorem 12.7.
[Improved version of Theorem A.1 of [GZ22b]] Suppose for some $\eta>0$, we have $k=o(p^{1/3-\eta})$. Denote $\alpha=(1+\eta)/(2+\eta)$. Then if $k\leq\sigma^{2}\leq 3k$, it holds that for every sufficiently large constant $D_{0}$ and $n\leq\frac{\alpha k\log(p/k)}{\log D_{0}}$, the cardinality of the set
@@
We will use three auxiliary lemmas from [GZ22b].
Lemma 12.8.
[Lemma A.5 of [GZ22b]] If $m_{1},m_{2}\in\mathbb{N}$ with $m_{1}\geq 4$ and $m_{2}\leq\sqrt{m_{1}}$, then
@@
Lemma 12.9.
[Lemma A.6 of [GZ22b]] The function $f:[0,1)\rightarrow\mathbb{R}$ defined by
@@
$\log p_{t,y}\geq\log t-t^{2}-y^{2}+\frac{1}{2}\log\frac{2}{\pi}.$ (12.4)
Lemma 12.10.
[Lemma A.7 of [GZ22b]] For any $t>0$, $y\in\mathbb{R}$, and $\rho\in[0,1)$,
@@
To prove Theorem 12.7, we note that for any $x\in\mathbb{R}^{n}$, $n^{-\frac{1}{2}}\lVert x\rVert_{2}\leq\lVert x\rVert_{\infty}$. It will therefore suffice to lower bound $Z_{s,\infty}$, where $Z_{s,\infty}$ is a counting random variable that is equal to the cardinality of the set
@@
$s=R\sqrt{k+\sigma^{2}}\exp\left(-\frac{k\log\frac{p}{k}}{n}\right)=\sqrt{k+\sigma^{2}}\exp\left(-(1-\alpha)\frac{k\log\frac{p}{k}}{n}\right)$ (12.5)
@@
where we define $t:=s/\sqrt{k}$. Proposition 12.11 concerns a truncated form of $\Upsilon$. Specifically, we show that $\mathbb{E}\left(\min\left\{1,\Upsilon-1\right\}\right)$ converges to 0 in expectation.
Proposition 12.11.
[New version of Proposition A.8 of [GZ22b]] For any sufficiently large constant $D>0$, if $n\leq\frac{k\log\frac{p}{k^{2}}}{2\log D}$ and $k=o(p^{1/3-\eta})$ for some $\eta>0$, then for $s$ per (12.5) and $t:=s/\sqrt{k}$, we have
@@
Proof of Proposition 12.11.
Throughout the proof of Proposition 12.11, we will take
@@
$t=s/\sqrt{k}=\sqrt{1+\frac{\sigma^{2}}{k}}\exp\left(-(1-\alpha)\frac{k\log\frac{p}{k}}{n}\right).$
As shown in section A.5 of [GZ22b], $\Upsilon$ can equivalently be written as
$\Upsilon=\Upsilon(Y)=\sum_{\ell=0}^{k}\frac{\binom{p}{k-\ell,k-\ell,\ell,p-2k-\ell}}{\binom{p}{k}^{2}}\prod_{i=1}^{n}\frac{q_{t,\frac{Y_{i}}{\sqrt{k}},\frac{\ell}{k}}}{p_{t,\frac{Y_{i}}{\sqrt{k}}}^{2}}$ (12.6)
@@
$\binom{k}{\ell}\leq\left(\frac{ke}{\ell}\right)^{\ell}$ (12.7)
@@
$\binom{p-k}{k-\ell}\leq\frac{\left(p-k\right)^{k-\ell}}{\left(k-\ell\right)!}$ (12.8)
Our assumption on $k$ implies $k\leq\sqrt{p}$, so we can apply Lemma 12.8 to get
$\binom{p}{k}\geq\frac{p^{k}}{4k!}$ (12.9)
We combine (12.7), (12.8), and (12.9) to see that for every $1\leq\ell\leq k$, it holds that
@@
$\Upsilon\leq 1+4\sum_{\ell=1}^{k}\left(\frac{p\ell}{ek^{2}}\right)^{-\ell}\prod_{i=1}^{n}\frac{q_{t,\frac{Y_{i}}{\sqrt{k}},\frac{\ell}{k}}}{p_{t,\frac{Y_{i}}{\sqrt{k}}}^{2}}$ (12.10)
@@
Observe that if $\Upsilon\geq 1+\zeta^{n}$, then the upper bound on $\Upsilon$ in (12.10) implies that at least one of the summands of
@@
$\mathbb{E}\left(\min\left\{1,\Upsilon-1\right\}\right)\leq\zeta^{n}+\sum_{\rho=\frac{1}{k},\frac{2}{k},\dots,1}\mathbb{P}(\Upsilon_{\rho})$ (12.11)

### v3

@@ Appendix A Deferred Proofs
Theorem A.2.
[Theorem 5.2 of [6]] Let $n,k,p\in\mathbb{N}$ with $k\leq p$. Suppose $Z\in\mathbb{R}^{n\times p}$ has i.i.d. standard Gaussian entries. Then for every $\delta>0$, there exists a constant $C=C_{\delta}>0$ such that if $n\geq Ck\log\frac{p}{k}$, then $Z$ satisfies the $k$-RIP with restricted isometric constant $\delta$ w.h.p. as $k\rightarrow\infty$.
@@
Finally, we will need two results from [27] concerning D.L.M.s
Proposition A.5.
[Proposition F.7 of [27]] Let $n,p,k\in\mathbb{N}$ with $k\leq\frac{1}{3}p$. Suppose that $X\in\mathbb{R}^{n\times p}$ satisfies the $3k$-RIP with restricted isometry constant $\delta_{3k}<\frac{1}{12}$. Then there is no $\frac{1}{4}$-D.L.M. triplet $(a,b,c)$ with respect to the matrix $X$ with $\lVert a\rVert_{2}^{2}+\lVert b\rVert_{2}^{2}\geq\frac{1}{4}\lVert c\rVert_{2}^{2}$.
Lemma A.6.
[Proposition F.11 of [27]] There exists a universal constant $c_{0}>0$ such that for any fixed triplet $(a,b,c)$ with $a\neq 0$,
$\mathbb{P}\left((a,b,c)\text{ is a }\frac{1}{4}\text{-D.L.M. triplet}\right)\leq 2\exp\left(-c_{0}n\min\left\{1,\frac{\|a\|_{2}^{2}+\|b\|_{2}^{2}}{\|c\|_{2}^{2}}\right\}\right),$
where for the case $c=0$ we abuse the notation by defining $\frac{1}{0}:=+\infty$.
The proof of Lemma A.6 is identical to the proof of Lemma F.11 of [27], but with $\frac{1}{4}$ in place of $\frac{1}{2}$.
The next proposition is the main technical contribution of this section.
Proposition A.7.
[(Improved) Proposition F.8 of [27]] Let $n,p,k\in\mathbb{N}$ with $k\leq\frac{1}{3}p$. Suppose that $X\in\mathbb{R}^{n\times p}$ has i.i.d. $\mathcal{N}(0,1)$ entries. There exists a constant $C_{1}>0$ such that if $n\geq C_{1}k\log(p/k)$ then w.h.p., there is no $\frac{1}{4}$-D.L.M. triplet $(a,b,c)$ with respect to some sets $\emptyset\neq S_{1},S_{2},S_{3}\subset[p]$ and the matrix $X$ such that the following conditions are satisfied.
@@
    $\lVert a\rVert_{2}^{2}+\lVert b\rVert_{2}^{2}+\lVert c\rVert_{2}^{2}\leq 1.25k/\log p$
Proof of Proposition A.7.
We first choose $C_{1}>0$ large enough based on Theorem A.2 so that $n\geq C_{1}k\log(p/k)$ implies that $X$ satisfies the $3k$-RIP with $\delta_{3k}<\frac{1}{16}$ w.h.p. In particular, all the probability calculations below will be conditioned on this high-probability event. We define the following sets parametrized by $r>0$ and $\alpha\in(0,1)$
$B_{r}:=\{(a,b,c)\big{|}a,b,c\in\{-1,0,1\}^{p}\setminus\{(0,0,0\},\|a\|_{0}+\|b\|_{0}+\|c\|_{0}\leq\min\left\{2k+1,r^{2}\right\}\}$
and
[…]
We call a triplet of sets $\emptyset\neq S_{1},S_{2},S_{3}\subseteq[p]$ good if
  • $S_{1},S_{2},S_{3}$ are pair-wise disjoint
  • $|S_{1}|=|S_{2}|$, $p\in S_{3}$ and $S_{1}\cup S_{3}=[k]\cup\{p\}$
For $\alpha\in\mathbb{R}$ and $S\subseteq\mathbb{R}$ we define the set
$S-\alpha:=\{s-\alpha|s\in S\}.$
For $i=1,2,3$ we set $P_{i}:=\{(i-1)p+1,(i-1)p+2,\ldots,ip\}$. Notice that the sets $P_{1},P_{2},P_{3}$ partition $[3p]$. We define the following family of subsets of $[3p]$,
$\mathcal{T}:=\{T\subset[3p]|\text{ the triplet }T\cap P_{1},T\cap P_{2}-p,T\cap P_{3}-2p\text{ is good}\}.$
It is easy to see that $\mathcal{T}\subset\{T\subset[3p]||T|\leq 2k+1\}$. Furthermore for any $T\in\mathcal{T}$ we define
$B_{r}(T):=\{(a,b,c)\in B_{r}\big{|}\mathrm{Support}\left((a,b,c)\right)\subseteq T,T\cap P_{1}=\mathrm{Support}\left(a\right)\}$
and
[…]
$T\cap P_{1},T\cap P_{2}-p,T\cap P_{3}-2p\}.$
We claim that
$D_{\frac{1}{4},r}=\bigcup_{T\in\mathcal{T}}D_{\frac{1}{4},r}\left(T\right).$ (A.1)
For the one direction, if $A=(a,b,c)\in D_{\frac{1}{4},r}\left(T\right)$ for some $T\in\mathcal{T}$ then $(a,b,c)$ is $\alpha$-DLM with corresponding super-supports $T\cap P_{1},T\cap P_{2}-p,T\cap P_{3}-2p$ which can be easily checked to satisfy assumption (2) of Proposition A.7 based on our assumptions. For the other direction, if $A\in D_{\frac{1}{4},r}$ is an $\alpha$-DLM with respect to $S_{1},S_{2},S_{3}$ satisfying assumption (2) of Proposition A.7, it can be easily verified that for the set $T=S_{1}\cup\left(S_{2}+p\right)\cup\left(S_{3}+2p\right)$ it holds $T\in\mathcal{T}$ and furthermore $A\in D_{\frac{1}{4},r}\left(T\right)$.
Now to prove the proposition it suffices to prove that there exists a constant $C_{1}>0$ such that if $n\geq C_{1}k\log(p/k)$ and $r^{2}\leq 1.25k/\log p$ then
$\lim_{k\rightarrow+\infty}\mathbb{P}\left(D_{\frac{1}{4},r}\neq\emptyset\right)=0.$
Using (A.1) with $\alpha=\frac{1}{4}$ and a union bound, it suffices to be shown that for some $C_{1}>0$ if $n\geq C_{1}k\log(p/k)$ and $r^{2}\leq 1.25k/\log p$ then
$\lim_{k\rightarrow+\infty}\sum_{T\in\mathcal{T}}\mathbb{P}\left(D_{\frac{1}{4},r}\left(T\right)\neq\emptyset\right)=0.$
But by Markov’s inequality for all such $T\in\mathcal{T}$,
$\mathbb{P}\left(|D_{\frac{1}{4},r}|\geq 1\right)\leq\mathbb{E}{|D_{\frac{1}{4},r}|}.$ (A.2)
Furthermore, for all $T\in\mathcal{T}$, $1\leq|T\cap P_{2}|\leq k$. By (A.2) and summing over the possible values of $|T\cap P_{2}|$ for $T\in\mathcal{T}$, it therefore suffices to show that for some $C_{1}>0$, if $n\geq C_{1}k\log(p/k)$ and $r^{2}\leq 1.25k/\log p$ then,
$\lim_{k\rightarrow+\infty}\sum_{m=1}^{k}\sum_{T\in\mathcal{T},|T\cap P_{2}|=m}\mathbb{E}\left(|D_{\frac{1}{4},r}(T)|\right)=0$ (A.3)
Fix $m\in[k]$ and a set $T\in\mathcal{T}$ with $|T\cap P_{2}|=m$. Then for any $A=(a,b,c)\in D_{\frac{1}{4},r}(T)$, since $D_{\frac{1}{4},r}(T)\subseteq B_{r}(T)$, we have $\|a\|_{2}^{2}+\|b\|_{2}^{2}+\|c\|_{2}^{2}\leq r^{2}$. Based on the definition of $D_{\frac{1}{4},r}(T)$, we also have $|\mathrm{Support}(a)|=|S_{1}|=|S_{2}|=|T\cap P_{2}|=m$. Hence, $\|a\|_{2}^{2}=m$ (as $a\in\{-1,0,+1\}^{p}$) and $\|c\|_{2}^{2}\leq\|a\|_{2}^{2}+\|b\|_{2}^{2}+\|c\|_{2}^{2}\leq r^{2}$. By Lemma A.6 we know that for any triplet $A=(a,b,c)$,
$\mathbb{P}\left(A\in D_{\frac{1}{4},r}(T)\right)\leq\exp\left(-c_{0}n\min\left\{1,\frac{\|a\|_{2}^{2}+\|b\|_{2}^{2}}{\|c\|_{2}^{2}}\right\}\right)$
for some universal constant $c_{0}>0$. Hence using the above inequalities we can conclude that for any such $A=(a,b,c)$ it holds that
$\mathbb{P}\left(A\in D_{\frac{1}{4},r}(T)\right)\leq 2\exp\left(-c_{0}n\min\left\{1,\frac{m}{r^{2}}\right\}\right)$ (A.4)
Linearity of expectation, the above bound, and an easy bound on the cardinality of the vectors in $\{-1,0,1\}^{p}$ with non-zero values only in $T$ and sparsity level at most $r^{2}$ imply that
$\mathbb{E}{|D_{\frac{1}{4},r}(T)|}\leq 3^{r^{2}}\binom{2k+1}{r^{2}}\exp\left(-c_{0}n\min\left\{1,\frac{m}{r^{2}}\right\}\right).$ (A.5)
We now count the number of possible $T\in\mathcal{T}$ with $|T\cap P_{2}|=m$. Recall that any $T\subseteq[3p]$ satisfies $T\in\mathcal{T}$ if and only if the triplet of sets $T\cap P_{1},T\cap P_{2}-p,T\cap P_{3}-2p$ is a good triplet. That is if and only if
  • (1)
    $T\cap P_{1},T\cap P_{2}-p,T\cap P_{3}-2p$ are pairwise disjoint sets and $|T\cap P_{1}|=|T\cap P_{2}-p|=|T\cap P_{2}|=m$
  • (2)
    $p\in T\cap P_{3}-2p$
  • (3)
    $\left(T\cap P_{1}\right)\cup\left(T\cap P_{3}-2p\right)=[k]\cup\{p\}$
Since a set $T\subseteq[3p]$ is completely characterized by the intersections with $P_{1},P_{2},P_{3}$, it suffices to count the number of triplets of sets $T\cap P_{i}$, $i=1,2,3$ satisfying the three above conditions. Now conditions (1),(3) imply that $T\cap P_{3}$ is completely characterized by $T\cap P_{1}$. Furthermore by checking conditions $(1),(2),(3)$ we know that $T\cap P_{1}$ is an arbitrary subset of $[k]$ of size $m$. Hence we have $\binom{k}{m}$ choices for both the sets $T\cap P_{1}$ and $T\cap P_{3}$. Finally for the set $T\cap P_{2}$ we only have that it needs to satisfy $|T\cap P_{2}|=m$. Hence for $T\cap P_{2}$ we have $\binom{p}{m}$ choices, giving in total that the number of sets $T\in\mathcal{T}$ with $|T\cap P_{2}|=m$ equals to $\binom{k}{m}\binom{p}{m}$. Hence,
$\sum_{T\in\mathcal{T},|T\cap P_{2}|=m}\mathbb{E}\left(|D_{\frac{1}{4},r}(T)|\right)\leq\binom{k}{m}\binom{p}{m}3^{r^{2}}\binom{2k+1}{r^{2}}\exp\left(-c_{0}n\min\left\{1,\frac{m}{r^{2}}\right\}\right).$
Summing over all $m=1,2,\ldots,k$ and using the bounds $\binom{k}{m}\leq\binom{p}{m}\leq(pe/m)^{m},3^{r^{2}}\binom{2k+1}{r^{2}}\leq(10k)^{r^{2}}$ we conclude that
$\sum_{m=1}^{k}\sum_{T\in\mathcal{T},|T\cap P|=m}\mathbb{E}\left(|D_{\frac{1}{4},r}(T)|\right)$
is at most
$k\max_{m=1,\ldots,k}\left[(pe/m)^{2m}(10k)^{r^{2}}\exp\left(-c_{0}n\min\left\{1,\frac{m}{r^{2}}\right\}\right)\right].$
Therefore it suffices to show that for some $C_{1}>0$ if $n\geq C_{1}k\log(p/k)$, $r^{2}\leq 1.25k/\log p$ then
$\lim_{k\rightarrow\infty}k\max_{m=1,\ldots,k}\left[(pe/m)^{2m}(10k)^{r^{2}}\exp\left(-c_{0}n\min\left\{1,\frac{m}{r^{2}}\right\}\right)\right]=0.$ (A.6)
Since this is a decreasing quantity in $n$ we plug in $n=C_{1}k\log(p/k)$. After taking logarithms it suffices to be proven that for $C_{1}$ large enough but constant, if $r^{2}\leq 1.25k/\log p$ then
$\max_{m=1,\ldots,k}\left[2m\log(pe/m)+r^{2}\log\left(10k\right)-c_{0}C_{1}k\log(p/k)\min\left\{1,\frac{m}{r^{2}}\right\}\right]+\log k\rightarrow-\infty.$
We consider the two cases: when $m\leq r^{2}$ and when $m\geq r^{2}$. Suppose $m\geq r^{2}$, that is $\min\left\{1,\frac{m}{r^{2}}\right\}=1$. Since $r^{2}\leq 1.25k/\log(p)$ it therefore holds for large enough $k$ that
$\max_{k\geq m\geq r^{2}}\left[2m\log(pe/m)+r^{2}\log\left(10k\right)-c_{0}C_{1}k\log(p/k)\min\left\{1,\frac{m}{r^{2}}\right\}\right]+\log k$
$\leq 2k\log(pe/k)+2k-c_{0}C_{1}k\log(p/k)+\log k$
$\leq-(c_{0}C_{1}-5)k\log(p/k),$
which if $C_{1}>6/c_{0}$ clearly diverges to $\infty$- ∞ as $k\rightarrow+\infty$.
Now suppose $m\leq r^{2}$, that is when $\min\left\{1,\frac{m}{r^{2}}\right\}=\frac{m}{r^{2}}$. We have
$\max_{1\leq m\leq r^{2}}\left[2m\log(pe/m)+r^{2}\log\left(10k\right)-c_{0}C_{1}k\log(p/k)\min\left\{1,\frac{m}{r^{2}}\right\}\right]+\log k$
$\leq\max_{1\leq m\leq r^{2}}\left[2m\log(pe)+r^{2}\log\left(10k\right)-c_{0}C_{1}k\log p\left(\frac{m}{r^{2}}\right)\right]+\log k.$
We write
$2m\log(pe)+r^{2}\log\left(10k\right)-c_{0}C_{1}k\log(p/k)\frac{m}{r^{2}}$
$=2m\log(pe)-\frac{c_{0}C_{1}}{2}k\log(p/k)\cdot\frac{m}{r^{2}}+r^{2}\log\left(10k\right)-\frac{c_{0}C_{1}}{2}k\log(p/k)\cdot\frac{m}{r^{2}}.$
Since for large enough $k$, $r^{2}\leq 2k/(\log(pe))$, we therefore see that
$2m\log(pe)-\frac{c_{0}C_{1}}{2}k\log(p/k)\cdot\frac{m}{r^{2}}\leq\left(2-\frac{c_{0}C_{1}}{4}\log(p/k)\right)m\log(pe)\leq-2\log p$ (A.7)
for $C_{1}\geq 16/(c_{0}\log(p/k))$ (note that this is at most a constant because $k\leq p/3$). Now we will bound the second summand. Again assuming $C_{1}>6/c_{0}$ and using that $m\geq 1$ we have
$r^{2}\log\left(10k\right)-\frac{c_{0}C_{1}}{2}k\log(p/k)\cdot\frac{m}{r^{2}}$ $\leq r^{2}(2\log p-c_{0}C_{1}k/r^{2})/2$ (A.8)
$\leq-(c_{0}C_{1}-2)\log p/2.$ (A.9)
Now combining (A.7) and (A.8) we conclude that for large enough $C_{1}>2/c_{0}$ that
$\max_{1\leq m\leq r^{2}}\left[2m\log(pe)+r^{2}\log\left(10k\right)-c_{0}C_{1}k\log p\frac{m}{r^{2}}\right]+\log k$
$\leq-2\log p-(c_{0}C_{1}-2)\log p/2+\log k$
$\leq-\log p\rightarrow-\infty,\text{ as }n,p,k\rightarrow+\infty$
which completes the proof of Proposition A.7. ∎
Now we have everything we need to prove Lemma 8.4.
@@
Choose $C>0$ large enough so that $X^{\prime}$ satisfies $3k$-RIP (Restricted Isometry Property) with $\delta_{3k}<\frac{1}{12}$ (we know we can do this by Theorem A.2). Let $T$ denote the support of $v$, and without loss of generality, assume that the support of $v^{*}$ is $[k]$. Denote $m=|T\setminus[k]|=|[k]\setminus T|$.
@@
$\lVert Y-Xv^{\prime}\rVert^{2}_{2}\geq\lVert Y-Xv\rVert^{2}_{2}-\frac{1}{4}n$ (A.10)
@@
$\lVert Xv^{*}+W-Xv+X_{j}-X_{i}\rVert_{2}^{2}\geq\lVert Xv^{*}+W-Xv\rVert^{2}_{2}-\frac{1}{4}n$ (A.11)
@@
Observe that we can rewrite the inequality (A.11) above in terms of $X^{\prime}$ to get that for all $i\in[p]$ and for all $j\in T$, we have
$\lVert X^{\prime}a+X^{\prime}b+X^{\prime}c-a_{i}X^{\prime}_{i}-b_{j}X^{\prime}_{j}\rVert^{2}_{2}\geq\lVert X^{\prime}(a+b+c)\rVert_{2}^{2}-\frac{1}{4}n$ (A.12)
This therefore gives that $(a,b,c)$ is $\frac{1}{4}$-DLM with respect to $[k]\setminus T$, $T\setminus[k]$, $\{p+1\}$, and the matrix X’. Therefore since $X^{\prime}$ satisfies $3k$-RIP with $\delta_{3k}<\frac{1}{12}$, by Proposition A.5, we have $\lVert a\rVert_{2}^{2}+\lVert b\rVert_{2}^{2}\leq\frac{1}{4}\lVert c\rVert^{2}_{2}$, or equivalently
@@
We therefore apply Proposition A.7 to conclude that it must be the case that with high probability,
$\frac{1.25k}{\log p}<|\mathrm{supp}(a)|+|\mathrm{supp}(b)|+\sigma^{2}\leq\frac{5}{4}\sigma^{2}$
@@
$\frac{k}{\log p}<\sigma^{2}$
and we get a contradiction. ∎
A.2. Proof of Theorem 10.1
Here we prove Theorem 10.1. The strategy here is that we will first prove a similar result for a pure noise model (i.e. with no signal). We then show how the linear model can be reduced to the pure noise model. We again follow [26]. Note that, as we mentioned also above, the first part of Theorem 10.1 is unchanged from [26], so we only prove the second part.
A.2.1. The Pure Noise Model
@@
Here there is no “true” vector $v^{*}$; we are simply trying to find the exactly $k$-sparse binary vector that minimises this squared error. Denote the optimal value of this optimization problem by $\psi_{2}$. Theorem A.8 is our main result for this subsection. It improves Theorem A.1 of [27] by reducing the assumption that $\Omega(k\log k)=n=o(k\log p)$ (and hence $k=p^{o(1)}$) to $k=o(p^{1/3-\eta})$ for some $\eta>0$ and no assumption on $n$.
Theorem A.8.
[Improved version of Theorem A.1 of [27]] Suppose for some $\eta>0$, we have $k=o(p^{1/3-\eta})$. Denote $\alpha=(1+\eta)/(2+\eta)$. Then if $k\leq\sigma^{2}\leq 3k$, it holds that for every sufficiently large constant $D_{0}$ and $n\leq\frac{\alpha k\log(p/k)}{\log D_{0}}$, the cardinality of the set
@@
We will use three auxiliary lemmas from [27].
Lemma A.9.
[Lemma A.5 of [27]] If $m_{1},m_{2}\in\mathbb{N}$ with $m_{1}\geq 4$ and $m_{2}\leq\sqrt{m_{1}}$, then
@@
Lemma A.10.
[Lemma A.6 of [27]] The function $f:[0,1)\rightarrow\mathbb{R}$ defined by
@@
$\log p_{t,y}\geq\log t-t^{2}-y^{2}+\frac{1}{2}\log\frac{2}{\pi}.$ (A.13)
Lemma A.11.
[Lemma A.7 of [27]] For any $t>0$, $y\in\mathbb{R}$, and $\rho\in[0,1)$,
@@
To prove Theorem A.8, we note that for any $x\in\mathbb{R}^{n}$, $n^{-\frac{1}{2}}\lVert x\rVert_{2}\leq\lVert x\rVert_{\infty}$. It will therefore suffice to lower bound $Z_{s,\infty}$, where $Z_{s,\infty}$ is a counting random variable that is equal to the cardinality of the set
@@
$s=R\sqrt{k+\sigma^{2}}\exp\left(-\frac{k\log\frac{p}{k}}{n}\right)=\sqrt{k+\sigma^{2}}\exp\left(-(1-\alpha)\frac{k\log\frac{p}{k}}{n}\right)$ (A.14)
@@
where we define $t:=s/\sqrt{k}$. Proposition A.12 concerns a truncated form of $\Upsilon$. Specifically, we show that $\mathbb{E}\left(\min\left\{1,\Upsilon-1\right\}\right)$ converges to 0 in expectation.
Proposition A.12.
[New version of Proposition A.8 of [27]] For any sufficiently large constant $D>0$, if $n\leq\frac{k\log\frac{p}{k^{2}}}{2\log D}$ and $k=o(p^{1/3-\eta})$ for some $\eta>0$, then for $s$ per (A.14) and $t:=s/\sqrt{k}$, we have
@@
Proof of Proposition A.12.
Throughout the proof of Proposition A.12, we will take
@@
$t=s/\sqrt{k}=\sqrt{1+\frac{\sigma^{2}}{k}}\exp\left(-(1-\alpha)\frac{k\log\frac{p}{k}}{n}\right)$
As shown in section A.5 of [27], $\Upsilon$ can equivalently be written as
$\Upsilon=\Upsilon(Y)=\sum_{\ell=0}^{k}\frac{\binom{p}{k-\ell,k-\ell,\ell,p-2k-\ell}}{\binom{p}{k}^{2}}\prod_{i=1}^{n}\frac{q_{t,\frac{Y_{i}}{\sqrt{k}},\frac{\ell}{k}}}{p_{t,\frac{Y_{i}}{\sqrt{k}}}^{2}}$ (A.15)
@@
$\binom{k}{\ell}\leq\left(\frac{ke}{\ell}\right)^{\ell}$ (A.16)
@@
$\binom{p-k}{k-\ell}\leq\frac{\left(p-k\right)^{k-\ell}}{\left(k-\ell\right)!}$ (A.17)
Our assumption on $k$ implies $k\leq\sqrt{p}$, so we can apply Lemma A.9 to get
$\binom{p}{k}\geq\frac{p^{k}}{4k!}$ (A.18)
We combine (A.16), (A.17), and (A.18) to see that for every $1\leq\ell\leq k$, it holds that
@@
$\Upsilon\leq 1+4\sum_{\ell=1}^{k}\left(\frac{p\ell}{ek^{2}}\right)^{-\ell}\prod_{i=1}^{n}\frac{q_{t,\frac{Y_{i}}{\sqrt{k}},\frac{\ell}{k}}}{p_{t,\frac{Y_{i}}{\sqrt{k}}}^{2}}$ (A.19)
@@
Observe that if $\Upsilon\geq 1+\zeta^{n}$, then the upper bound on $\Upsilon$ in (A.19) implies that at least one of the summands of
@@
$\mathbb{E}\left(\min\left\{1,\Upsilon-1\right\}\right)\leq\zeta^{n}+\sum_{\rho=\frac{1}{k},\frac{2}{k},\dots,1}\mathbb{P}(\Upsilon_{\rho})$ (A.20)